# Distributional Evaluation of Generative Models via Relative Density Ratio

### v1

@@ 3.2.2 Feedforward Neural Network Formulation
This is a commonly used method of neural network parameterization in the literature (Suzuki and Nitanda, 2021; Kaji et al., 2023).
@@ 3.2.3 Convergence Rate for Sieved Neural Network
Denote $\mathcal{G}_{n}$ as the sieve space of neural networks. The sieved M-estimator allows the search space $\mathcal{G}_{n}$ to grow with sample size $n$, and is eventually flexible enough to approximate the true function (see details in Section 3.4 of van der Vaart and Wellner (1996)). We define the sieve space according to the neural network parameter scales in Proposition 2 in Suzuki and Nitanda (2021). Let $\mathcal{G}_{n}=\Phi(L_{N},U_{N},S_{N},B_{N})$, where the neural network parameters depend on $N$, and $N=N(n)$ depends on the sample size $n$. The dependence of $(L_{N},U_{N},S_{N},B_{N})$ on $N$ is specified in Proposition 2 in Suzuki and Nitanda (2021) and in our proof (25). The rate $N(n)$ will be specified later in the proof.
Now we can assume that the true density ratio $g_{0}\in B^{\beta}_{s,t}(\mathcal{X})$. Denote $\tilde{\beta}=\left(\sum_{j=1}^{d}1/\beta_{j}\right)^{-1}$ as the harmonic mean of the smoothness vector $\beta$. Let $\mathcal{U}\left(B^{\beta}_{s,t}(\mathcal{X})\right)$ be the unit ball of $B^{\beta}_{s,t}(\mathcal{X})$.
@@
Previously, in the proof of Theorem 3, consistency is required; the true function $g_{0}$ must be in the same function space $\mathcal{G}$ as the estimator; and the entropy condition is prespecified in Assumptions 4 and 6. Here in Theorem 4, consistency is implied but not required, and we allow $g_{0}$ to be outside of $\mathcal{G}_{n}$, although the distance between $\mathcal{G}_{n}$ and $\mathcal{U}\left(B^{\beta}_{s,t}(\mathcal{X})\right)$ goes to 0 as $N\to\infty$ (see Proposition 2 of Suzuki and Nitanda (2021)); the entropy condition of the sparse neural network space $\mathcal{G}_{n}$ is shown in Lemma 4 rather than being prespecified. Similar to Corollary 1, since the proof of Theorem 4 does not require the independence of $P_{n}$ and $Q_{m}$, we can directly apply this result to the general relative density ratio $r(x)$ by replacing $Q$ with $\widetilde{Q}=\alpha P+(1-\alpha)Q$, and similarly for $\widetilde{Q}_{m}$.
@@
Remark. The convergence rate in (13) depends only on the smoothness parameter $\tilde{\beta}$ and does not depend on the dimension $d$ of the sample space. This is because we allow the true function $g_{0}$ to have different smoothness in different directions, and the harmonic mean $\tilde{\beta}$ is only driven by a few least smooth directions, rather than assuming a spatially homogeneous smoothness bound in all directions (Hölder space), where the dimension $d$ will come into play, as in previous works (Nakada and Imaizumi, 2020).
@@ 4 Numerical Experiments
We use three numerical experiments to illustrate the rich information that can be extracted from the relative density ratio $r(x)$: the MNIST dataset (LeCun et al., 1998), the CelebA-64 dataset (Liu et al., 2015), and the microbiome data from the American Gut Project (McDonald et al., 2018b). Different from Simon et al. (2019), where the evaluation is conducted on the extracted inception features rather than the raw data, we perform all of our numerical experiments on the raw data to preserve the authentic distributional discrepancies.
@@ 4.1 MNIST
The MNIST dataset (LeCun et al., 1998) contains 60,00060{,}000 training and 10,00010{,}000 testing grayscale images, and each image, of dimensions $28\times 28$, is preprocessed so that all the voxels are mapped to [1,1][-1,1]. We test on two pretrained generators with saved checkpoints provided in the GitHub repository (Sharma, 2019): VAE (Kingma and Welling, 2014) and DCGAN (Radford et al., 2015). To train the ratio function, we use batch updates and stochastic gradient descent to iterate through the training set; the architecture for the ratio function is the same as the discriminator in DCGAN, a CNN-based architecture, except for the output layer. To ensure the theoretical bound $[0,2]$ for $r(x)$, we apply a Bounded Softplus output layer: map the Softplus ($\log(1+\exp(x))$) output of $0,\infty)$ to $[0,2]$ through $\frac{2x}{x+1}$. The ratio function is estimated using the training data, and we evaluate the ratio function on the 10,00010{,}000 real testing data and another 10,00010{,}000 generated samples from each generator.
@@
Figure 3 provides visualizations on $\{r(X_{p}),r(X_{q})\}$ for VAE, and Figure 4 for DCGAN. Immediately from the scalar metric, $h^{2}(p,\frac{p+q}{2})$ is 0.286 for VAE, and 0.162 for DCGAN, we know that DCGAN is closer to the real sample than VAE. Note that this distance $h^{2}(p,\frac{p+q}{2})$ directly results from the negative training loss and does not require separate computation since the variational form is the dual problem of the $\phi$-divergence. The sample-level RDR allows us to dive much deeper than simply picking the best fit.
@@
Figure 3: MNIST: VAE result. A: Histogram of $r(X_{p}),r(X_{q})$ evaluated for the real test data and generated images. B: Violin plot of $r(X_{p})$ stratified by digits label in the test data. C: Top row: Real test images; Bottom row: Fake generated images; From left to right: images with $r(X)$ values from the smallest, closest to 1, and the largest; the range of $r(X)$ for the shown images is in each subplot title.
@@
Panel A in Figures 3 and 4 shows the histogram of $\{r(X_{p}),r(X_{q})\}$. The majority of the VAE $r(X_{q})$ mass concentrates near 0 and has almost no overlap with $r(X_{p})$, indicating low coverage and low fidelity; whereas for DCGAN, the empirical support of $r(X_{q})$ is quite similar to that of $r(X_{p})$, indicating good coverage. But for DCGAN, there is still a noticeable portion of $r(X_{q})$ centered around 0, even though this portion is still in or near the mass of certain $r(X_{p})$ samples, indicating a portion of the real sample is overly represented by the generator DCGAN. The sample-level RDR enables us to examine which regions of the support the DCGAN is biased toward.
Figure 4: MNIST: DCGAN result. See explanations for the three groups of plots A, B, and C in the caption of Figure 3.
Because the real sample comes with digit labels, we can first check $r(X_{p})$ stratified by digits, as shown in Panel B of Figures 3 and 4. For VAE, although most real samples are not well represented, we notice that $r(X_{p})$ of digit 1 is comparably smaller than that of other digits, indicating that VAE can sample digit 1 slightly better than the other digits. Note that for better visualization, the y-axis limits differ in Figures 3 and 4. For DCGAN, Figure 4 B shows that , although $r(X_{p})$ for DCGAN ranges from 0 to nearly 1.9 for all digits, digits 1 and 6 appear to have the most mass above 1, which means DCGAN tends to under-represent digits 1 and 6.
@@
Panel C in Figures 3 and 4 provides instances of real and fake images whose $r(X)$ values are the closest to 0, 1, and 2, respectively. For VAE, because the ranges of both $r(X_{p})$ and $r(X_{q})$ are very narrow and non-overlapping, the left and middle panels are the same for the real sample (top row); the middle and right panels have overlapping instances in the fake sample (bottom row). We can still see that the left panel for the fake sample shows examples of poor fidelity, while the images in the right panel look more realistic. For the DCGAN result, the fake images are all realistic; however, but we can still check the over-represented (left panel) samples that contain many O-looking shapes such as 0,8, and 3, as well as the under-represented (right panel) samples, such as digit 1.

### v2

@@ 3.2.2 Feedforward Neural Network Formulation
This is a commonly used method of neural network parameterization in the literature (suzuki2021deep; kaji2023adversarial).
@@ 3.2.3 Convergence Rate for Sieved Neural Network
Denote $\mathcal{G}_{n}$ as the sieve space of neural networks. The sieved M-estimator allows the search space $\mathcal{G}_{n}$ to grow with sample size $n$, and is eventually flexible enough to approximate the true function (see details in Section 3.4 of van1996weak). We define the sieve space according to the neural network parameter scales in Proposition 2 in suzuki2021deep. Let $\mathcal{G}_{n}=\Phi(L_{N},U_{N},S_{N},B_{N})$, where the neural network parameters depend on $N$, and $N=N(n)$ depends on the sample size $n$. The dependence of $(L_{N},U_{N},S_{N},B_{N})$ on $N$ is specified in Proposition 2 in suzuki2021deep and in our proof (25). The rate $N(n)$ will be specified later in the proof.
Let $\mathcal{U}\left(B^{\beta}_{s,t}(\mathcal{X})\right)$ be the unit ball of $B^{\beta}_{s,t}(\mathcal{X})$. Denote $\tilde{\beta}=\left(\sum_{j=1}^{d}1/\beta_{j}\right)^{-1}$ as the harmonic mean of the smoothness vector $\beta$.
@@
Previously, in the proof of Theorem 3, consistency is required; the true function $g_{0}$ must be in the same function space $\mathcal{G}$ as the estimator; and the entropy condition is prespecified in Assumptions 4 and 6. Here in Theorem 4, consistency is implied but not required, and we allow $g_{0}$ to be outside of $\mathcal{G}_{n}$, although the distance between $\mathcal{G}_{n}$ and $\mathcal{U}\left(B^{\beta}_{s,t}(\mathcal{X})\right)$ goes to 0 as $N\to\infty$ (see Proposition 2 of suzuki2021deep); the entropy condition of the sparse neural network space $\mathcal{G}_{n}$ is shown in Lemma 4 rather than being prespecified. Similar to Corollary 1, since the proof of Theorem 4 does not require the independence of $P_{n}$ and $Q_{m}$, we can directly apply this result to the general relative density ratio $r(x)$ by replacing $Q$ with $\widetilde{Q}=\alpha P+(1-\alpha)Q$, and similarly for $\widetilde{Q}_{m}$.
@@
Remark. The convergence rate in (13) depends only on the smoothness parameter $\tilde{\beta}$ and does not directly depend on the dimension $d$ of the sample space. This is because we allow the true function $g_{0}$ to have different smoothness in different directions, and the harmonic mean $\tilde{\beta}$ is only driven by a few least smooth directions, rather than assuming a spatially homogeneous smoothness bound in all directions (Hölder space), where the dimension $d$ will come into play, as in previous works (nakada2020adaptive).
@@ 4 Numerical Experiments
We use three numerical experiments to illustrate the rich information that can be extracted from the relative density ratio $r(x)$: the MNIST dataset (mnist), the CelebA-64 dataset (CelebA64), and the microbiome data from the American Gut Project (AmericanGut). Different from pmlr-v97-simon19a, where the evaluation is conducted on the extracted inception features rather than the raw data, we perform all of our numerical experiments on the raw data to preserve the authentic distributional discrepancies. We provide simulation studies in Section F where the true relative density ratio function is known, and we compare the 4-layer MLP based on the Hellinger loss with other existing density ratio estimation methods such as uLSIF, KLIEP, and the density ratio trick as discussed in Section 2.3. We use the bounded sigmoid function $\sigma_{\alpha}(x)=\frac{2}{1-\exp\left\{-\alpha x\right\}}$ with $\alpha=1$ as the default output activation function, and add a discussion and sensitivity analysis on the difference choice of output activation function in the Supplementary Section D.
@@ 4.1 MNIST
The MNIST dataset (mnist) consists of 60,00060{,}000 training and 10,00010{,}000 testing grayscale images, and each image, with dimensions $28\times 28$, is preprocessed so that all the voxels are mapped to [1,1][-1,1]. We test on two pretrained generators with saved checkpoints provided in the GitHub repository (csinva2019ganvae): VAE (kingma2014vae) and DCGAN (DCGAN). To train the ratio function, we use batch updates and stochastic gradient descent to iterate through the training set; the architecture for the ratio function is the same as the discriminator in DCGAN, a CNN-based architecture. The ratio function is estimated using the training data, and we evaluate the ratio function on the 10,00010{,}000 real testing data and another 10,00010{,}000 generated samples from each generator.
@@
Figure 3 provides visualizations on $\{r(X_{p}),r(X_{q})\}$ for VAE, and Figure 4 for DCGAN. Immediately from the scalar metric, $h^{2}(p,\frac{p+q}{2})$ is 0.277 for VAE, and 0.135 for DCGAN, we know that DCGAN is closer to the real sample than VAE. Note that this distance $h^{2}(p,\frac{p+q}{2})$ directly results from the negative training loss and does not require separate computation since the variational form is the dual problem of the $\phi$-divergence. The sample-level RDR allows us to dive much deeper than simply picking the best generative model.
@@
Figure 3: MNIST: VAE result. A: Histogram of $r(X_{p}),r(X_{q})$ evaluated for the real test data and generated images. B: Top row: Real test images; Bottom row: Fake generated images; From left to right: images with $r(X)$ values from the smallest, closest to 1, and the largest; the range of $r(X)$ for the shown images is in each subplot title.
@@
Panel A in Figures 3 and 4 shows the histogram of $\{r(X_{p}),r(X_{q})\}$. The majority of the VAE $r(X_{q})$ mass concentrates near 0 and has almost no overlap with $r(X_{p})$, indicating low coverage and low fidelity. For DCGAN, the empirical support of $r(X_{q})$ overlaps with $r(X_{p})$ on $[0,2]$, indicating good coverage. But for DCGAN, there is still a noticeable portion of $r(X_{q})$ centered around 0, even though this portion still overlaps with the mass of certain $r(X_{p})$ samples, indicating a portion of the real sample is overly represented by the generator DCGAN. The sample-level RDR enables us to examine which regions of the support the DCGAN is biased toward. Because the real sample comes with digit labels, we can first check $r(X_{p})$ stratified by digits, as shown in Panel B of Figures 4. For DCGAN, although $r(X_{p})$ ranges from 0 to 2 for all digits, digits 1 and 6 appear to have the most mass above 1, which means DCGAN tends to under-represent digits 1 and 6.
Figure 4: MNIST: DCGAN result. See explanations for the three groups of plots A, B, and C in the caption of Figure 3. A: Histogram of $r(X_{p}),r(X_{q})$ evaluated for the real test data and generated images. B: Violin plot of $r(X_{p})$ stratified by digits label in the test data. C: Top row: Real test images; Bottom row: Fake generated images; From left to right: images with $r(X)$ values from the smallest, closest to 1, and the largest; the range of $r(X)$ for the shown images is in each subplot title.
@@
Figure 3(B) and Figure 4(C) provides examples of real and fake images whose values $r(X)$ are closest to 0, 1, and 2, respectively. For VAE, because the ranges of both $r(X_{p})$ and $r(X_{q})$ are very narrow and non-overlapping, the left and middle panels are the same for the real sample (top row); the middle and right panels have overlapping instances in the fake sample (bottom row). Although the generation quality of VAE is not ideal, it can still produce a few realistic images. The fake VAE samples on the left panel (near 0) illustrate poor fidelity, whereas those on the right appear more realistic (near 2). For the DCGAN result in Figure 4, the range of $r(X_{p})$ and $r(X_{q})$ almost fully overlaps. However, the RDR values indicate DCGAN has its own distributional bias compared to the real sample. The over-represented samples (left panel) in Figure 4(C) exhibit more rounded, spatially diffused patterns, such as “5” and “8”, across the 28×28 grid, whereas the under-represented samples (right panel) are either geometrically simpler such as digit “1”, or more irregular-looking.